Jill is 6 years older than Fernando. In two years, she will be twice as old as Fernando. How old is Jill now?
Answer by Alan3354(69443)

Jill is 6 years older than Fernando. In two years, she will be twice as old as Fernando. How old is Jill now?
---------
the difference is 6 years.
--> in 2 years 6 & 12
======================
The hard way:
J = F + 6
J+2 = 2*(F+8) = 2F + 16
-----
Sub for J
F+6+2 = 2F + 16
F+8 = 2F + 16
F = 8
J = 14